# Quasi-convex Splittings of Acylindrical Graphs of Locally finite-height Groups

### v1

@@ Abstract.
We find a condition on the action of a finitely presented group on a simplicial tree which guarantees that this action will be dominated by an acylindrical action with finitely generated edge stabilisers, and find the first example of an action of a finitely presented group where there is no such dominating action. As a consequence, we show that any hyperbolic group that admits a decomposition as an acylindrical graph of (possibly infinitely generated) free groups is virtually cocompact special, and that if one assumes that every finite-height subgroup of a hyperbolic group is quasi-convex then every finitely generated subgroup of a one-relator group with an acylindrical Magnus hierarchy admits a quasi-convex hierarchy.
@@ 1. Introduction
Let $G$ be a one-relator group. One of the oldest and most effective tools for studying such groups is a Magnus Hierarchy, an iterative HNN-structure on $G$ whose vertex groups are one-relator groups and whose edges are free subgroups generated by subsets of the generators of the vertices (c.f. 3.6). In his paper [Linton22], Linton proved that if a Magnus hierarchy for $G$ is acylindrical (c.f. 1.1) then $G$ is hyperbolic and the hierarchy is quasi-convex (c.f. 2.15), a much stronger condition.
Now assume that $G$ has such an acylindrical hierarchy, and let $H$ be a finitely generated subgroup of $G$. By coherence of $G$ [Coherence1r, Theorem 1.1] and Gersten’s theorem [Gerston, Theorem 5.4] (see [GerstenQ] for the case when $G$ has torsion) the group $H$ is hyperbolic, so a natural question to ask is whether $H$ will itself admit a quasi-convex hierarchy. Indeed, $H$ will always admit an acylindrical hierarchy induced by the hierarchy for $G$, but this need not be quasi-convex, the most immediate issue being that we cannot assume that it has finitely generated edge stabilisers — see Section 3.3 for an example of such a $G$ and $H$ where the induced hierarchy on $H$ has infinitely generated edge stabilisers. In this paper we prove that even in such a case, we may construct a quasi-convex hierarchy for $H$.
The definition of an acylindrical action on a tree was first formulated by Sela [Sela97], and later generalised by Weidmann [Weidmann07] to the following.
@@
We say that a group $G$ is acylindrically arboreal if $G$ admits an acylindrical action on a tree that is non-elementary (c.f. Theorem 2.6). Acylindrical arboreality is a special case of acylindrical hyperbolicity, a well studied and powerful generalisation of hyperbolicity that has been of much interest in recent years [minasyanOsin13, Osin13, Osin17], and was the first such generalisation to encompass mapping class groups of hyperbolic surfaces ([Bowditch08, MasurMinsky99], see also [Osin13, Section 8]).
A concept strongly related to an acylindrical action on a tree is that of a quasi-convex splitting, an action of a hyperbolic group $G$ on a simplicial tree where the edge stabilisers are all quasi-convex in $G$ (c.f Definition 2.15). A quasi-convex splitting with finitely many orbits of edges will always be acylindrical by a result of Gitik, Mj, Rips and Sageev [GitikMahanRipsSageev98, Main Theorem] that quasi-convex subgroups have finite-height (c.f. Definition 2.8), meaning that in an action on a tree with quasi-convex edge groups any large enough set of edges will have a finite stabiliser. The converse is more difficult, and is the subject of the following conjecture that can be found in question form in [BestvinaList].
@@
Let $G$ be a hyperbolic group, and $H$ a finitely generated subgroup of $G$ with finite-height. Then $H$ is quasi-convex in $G$.
@@
[Cohen23, Question 1.7] Let $G$ be a hyperbolic group with a non-elementary acylindrical action on a tree. Does $G$ admit a quasi-convex splitting?
One of the main reasons for the significance of quasi-convex splittings is their link to virtual specialness of hyperbolic groups. In particular, in [WiseQCBook, Theorem 13.1], Wise proves that if a hyperbolic group admits a quasi-convex splitting with virtually cocompact special vertex stabilisers then it itself is virtually cocompact special. Virtual specialness is an incredibly important concept in modern geometric group theory, and its definition along with the fact that it is inherited from vertex groups in a quasi-convex splittings was instrumental in proving the Virtual Haken Conjecture, one of the most influential group theoretic results in recent years, see [VirtualHakenAgol].
@@
    The fundamental groups of closed and orientable hyperbolic $3$-manifolds [Cohen23, Theorem 1.8]; and
@@
    Free groups, and more generally hyperbolic locally quasi-convex groups such as hyperbolic limit groups [Wilton2006HallsTF, Corollary 3.12] and many small cancellation groups [MccamondWiseLocalQC, MccamondWiseLocalQC2].
As stated above, the first obstacle to overcome when attempting to construct quasi-convex splittings from acylindrical ones is that a generic acylindrical splitting need not have finitely generated stabilisers. In [Cohen23], this problem was solved for hyperbolic $3$-manifold groups using a construction of Stallings, Epstein and Waldhausen which takes a splitting of such a group and outputs a collection of embedded surfaces one may cut along to obtain a new dominating action on a tree. This construction was proved to preserve acylindricity (and indeed promote it to quasi-convexity) in [Cohen23] using the Subgroup Tameness Theorem for hyperbolic $3$-manifolds (see [AschenbrennerFriedlWilton12, Theorem 5.2], for example).
A similar but weaker method exists in general for finitely presented groups. In particular, there is the construction due to Dunwoody known as the Dunwoody resolution [DunwoodyResolution, Lemma 2.2] that takes as input an action of a finitely presented group on a tree and outputs a dominating action with finitely generated edge stabilisers that are in some sense geometric, in that they arise as one-dimensional subspaces of a Cayley complex. For a locally quasi-convex group, the actions produced by the Dunwoody resolution will always have quasi-convex edge stabilisers, answering Question 1.3 in this case.
The main technical result of this paper exhibits a large class of actions in which acylindricity is preserved by the Dunwoody Resolution, and indeed where the resolution produces finite-height splittings that dominate the original acylindrical splitting. We say that a group $G$ is locally finite-height if every finitely generated subgroup of $G$ has finite-height in $G$, and that an action of $G$ on a simplicial tree $T$ is finite-height if every edge stabiliser of this action has finite height in $G$. The following is our main theorem.
@@
    If all vertex stabilisers of the action of $G$ on $T$ are locally finite-height then $G$ admits a non-elementary finite-height splitting with finitely generated edge stabilisers that dominates the action of $G$ on $T$. Thus, assuming Conjecture 1.2, if $G$ is hyperbolic then $G$ admits a quasi-convex splitting that dominates the action of $G$ on $T$.
@@
    If there exists an edge stabiliser that is finite-height in $G$, and this edge stabiliser is locally finite-height, then $G$ admits a non-elementary finite-height splitting with finitely generated edge stabilisers. Furthermore, if every edge stabiliser satisfies this condition then we may assume that the new action dominates the action of $G$ on $T$. Thus, assuming Conjecture 1.2, if $G$ is hyperbolic then $G$ admits a quasi-convex splitting that dominates the action of $G$ on $T$.
Theorem 1.4 is wide ranging — the main application we present here uses that all countable free groups are locally finite-height, and thus every finitely presented group that is an acylindrical graph of free groups will admit a finite height splitting with finitely generated edge stabilisers. Indeed, finitely generated free groups are known to be locally quasi-convex by Hall’s Theorem [Hall_1949], and so when dealing with hyperbolic graphs of free groups one may often obtain quasi-convexity without having to appeal to Conjecture 1.2. The following corollary follows immediately from Theorem 1.4(a) and an application of the Quasi-Convex Combination Theorem of Kapovich [Kapovich01, Theorem 1.2], which states that if a torsion free hyperbolic group $G$ acts acylindrically on a tree with each vertex stabiliser hyperbolic and each edge stabiliser quasi-convex in its neighbouring vertex stabilisers, then each edge stabiliser will be quasi-convex in $G$.
@@
Let $G$ be a hyperbolic group, and assume that $G$ acts acylindrically on some tree $T$ such that the vertex stabilisers of this action are free. Then $G$ admits a quasi-convex splitting with free vertex groups, and so $G$ is virtually cocompact special by [WiseQCBook, Theorem 13.1].
We are also able to use part (b) of Theorem 1.4 to prove the following, which settles our opening question up to Conjecture 1.2.
@@
Let $H$ be a finitely generated subgroup of a one-relator group $G$ such that $G$ admits an acylindrical Magnus hierarchy $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$. Then $H$ admits a non-trivial hierarchy of non-trivial finite-height splittings with finitely generated edge stabilisers. In particular the group $H$ is acylindrically arboreal if and only if it admits a non-trivial finite-height splitting with finitely generated edge stabilisers, which will be a quasi-convex splitting if Conjecture 1.2 is assumed.
Furthermore, if we assume Conjecture 1.2, then $H$ admits a quasi-convex hierarchy.
@@
A positive answer to Question 1.7 would have strong consequences. In particular, again using [WiseQCBook, Theorem 13.1], a positive answer would imply that every hyperbolic group that admits an acylindrical hierarchy in fact admits a quasi-convex hierarchy, and therefore is virtually cocompact special with a proof analogous to that of Corollary 1.6. This would provide a significantly weaker condition for virtual specialness of a hyperbolic group than we currently have at our disposal.
We finish by illustrating the issues that arrise when attempting to relax the hypotheses of Theorem 1.4. In Section 3.4 we will show that the Dunwoody resolution cannot be assumed to preserve the acylindricity of a general action with the following proposition.
@@
In particular, the Dunwoody resolution will not preserve acylindricity in general.
Whist this result is strong, the group constructed does admit an acylindrical action with finitely generated edge stabilisers. To counter this, we ask the following.
Question 1.9.
@@ 1.1. Acknowledgements
This work was completed while the author was a PhD student at the University of Cambridge, supervised by Jack Button, and the author is very grateful to his supervisor for all of his help. The author would also like to thank Henry Wilton for several very helpful discussions, and Daniel Groves for his helpful suggestions. Finally, financial support from the Cambridge Trust Basil Howard Research Graduate Studentship is gratefully acknowledged.
@@ 2. Preliminaries - Acylindrical Hyperbolicity and Acylindrical Arboreality
We will assume the reader has some familiarity with Bass–Serre theory, and for a more detailed discussion we refer to [stilwell2002trees, dicks2011groups]. We define a graph of groups $(\Gamma,\mathfrak{G})$ to be a connected directed graph $\Gamma$ that may not be finite or simple along with $\mathfrak{G}$, which is the following data:
@@
We define the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ of $(\Gamma,\mathfrak{G})$ as in [stilwell2002trees, Section I.5.1].
We will use a slight abuse of notation to consider each vertex group $G_{v}$ as a subgroup of the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ along the natural inclusion. Similarly, we will consider each edge group $G_{e}$ to be the subgroup of the fundamental group given by the image of $d_{0}(G_{e})$ in the vertex group $G_{i(e)}$. We call a graph of groups trivial if there exists some $v\in V(\Gamma)$ such that $G_{v}=\pi_{1}(\Gamma,\mathfrak{G})$, or non-trivial otherwise. We say that a graph of groups $(\Gamma,\mathfrak{G})$ is a graph of groups decomposition or splitting of a group $G$ if the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ is isomorphic to $G$, and identify $G$ with $\pi_{1}(\Gamma,\mathfrak{G})$. We denote by $T(\Gamma,\mathfrak{G})$ the Bass–Serre tree associated to the splitting, on which $G$ acts naturally by isometry with respect to the edge metric and without inversion as in [stilwell2002trees, Section I.5.3].
From now on we will assume that all trees are simplicial, and all actions on trees are by simplicial isometry and without inversion, so that every action on a tree is associated to a quotient graph of groups as in [stilwell2002trees, Section I.5], whose edge and vertex groups correspond to conjugacy classes of stabilisers of the original group action. We say that an action of a group $G$ on a tree $T$ is minimal if there exists no proper subtree of $T$ that is fixed setwise by the action of $G$. We have the following lemma, which will allow us to assume that the actions we consider are minimal in most cases.
@@
[dicks2011groups, Lemma I.4.13] Let $G$ be a finitely generated group acting on a tree $T$. Then there exists a unique minimal subtree $T_{m}\subseteq T$, and the induced action of $G$ on $T_{m}$ has finitely many orbits of edges.
The following lemma of Bieri will be essential in dealing with finite generation of stabilisers of action on trees.
@@
[bieri1981homological, Proposition 2.13(a) and (b)] Let $G$ be a finitely generated group acting on a tree $T$ with finitely many orbits of edges. If every edge stabiliser of this action is finitely generated, then so is every vertex stabiliser.
@@
    For a more complicated example, a right angled Artin group is acylindrically arboreal if and only if the underlying graph has graph theoretical diameter at least three [Cohen23, Corollary 3.12].
@@
    If $G$ is the fundamental group of a closed non-geometric $3$-manifold, then Wilton and Zalesskii prove that $G$ acts acylindrically on the tree corresponding to its JSJ-decomposition [WiltonZalesskii08, Lemma 2.4]. In the hyperbolic case, the fundamental group $G$ of a finite volume orientable hyperbolic $3$-manifold will be acylindrically arboreal if and only if the manifold contains an embedded $2$-sided closed incompressible surface whose fundamental group includes into $G$ as a geometrically finite subgroup [Cohen23, Theorem 1.6].
By a result first explicitly proved in [Cohen23, Theorem 2.17], but essentially due to Minasyan and Osin [minasyanOsin13, Lemma 4.2], an action on a simplicial tree is non-elementary acylindrical as in Definition 2.4 if and only if it is non-elementary acylindrical in the sense of Bowditch [Bowditch08]. Thus, if a group is acylindrically arboreal then it is acylindrically hyperbolic, and we have the following powerful classification theorem.
@@
[Osin13, Theorem 1.1] Let $G$ be a group acting acylindrically on a tree $T$. Then $G$ satisfies exactly one of the following conditions.
@@
[Osin13, Corollary 1.5] Let $G$ be an acylindrically arboreal group. Then every infinite normal subgroup of $G$ is acylindrically arboreal. In particular, such a subgroup cannot be abelian.
2.1. Finite-height and Quasi-Convex Splittings
As in Question 1.3, we are concerned with whether we can use acylindrical splittings to build splittings with stronger properties. In particular, we are concerned with finite-height and quasi-convex splittings of hyperbolic groups.
Intuitively, a finite-height splitting is an action on a tree where a set of edges containing enough elements has a finite stabiliser. It is important to note that, unlike acylindricity, this definition does not require that these stabilisers are uniformly bounded, only that they are finite. More formally, we define finite height as follows.
@@ 2.1. Finite-height and Quasi-Convex Splittings
Let $G$ be a group, and $H\leq G$ a subgroup. For $g_{1},g_{2}\in G$, we say that the conjugates $g_{1}Hg_{1}^{-1}$ and $g_{2}Hg_{2}^{-1}$ are essentially disjoint if $g_{1}H\neq g_{2}H$.
For some natural number $n$, we say that $H$ has height at most n in $G$ if the intersection of any collection of at least $n$ essentially disjoint conjugates of $H$ in $G$ is finite. We define the height of $H$, $h_{G}(H)$ to be the smallest $m$ such that $H$ has height at most $m$ in $G$, or to be infinite if no such $m$ exists.
@@
[…]
We say that the action of a group $G$ on a tree $T$ is a finite-height splitting if for all edges $e\in E(T)$, the $G$-stabiliser $G_{e}$ of $e$ is finite-height in $G$, i.e $h_{G}(G_{e})$ is finite. As above, some care should be taken when comparing acylindricity to finite height splittings owing to the different restrictions on the finite stabilisers. In particular, while in an acylindrical action stabilisers of long enough paths must be uniformly bounded, in a finite height splitting we only require that stabilisers of long enough sets are finite. The following example demonstrates this difference.
@@
The group $G$ acts on $T$ via the left coset action on the labels of its vertices, and the edge stabilisers of this action are all isomorphic to finite groups, so naturally have height $1$. This action is therefore finite-height, but will not be acylindrical, as for all $k$ we can find a path of length $k$ whose stabiliser is contains $2^{i}$ elements for any $i$ by taking the path that starts at the vertex labelled $(\mathbb{Z}/2^{i}\mathbb{Z})$ and finishing at the vertex labelled $(\mathbb{Z}/2^{i+k}\mathbb{Z})$.

### v2

# Quasi-convex Splittings of Acylindrical Graphs of Locally finite height Groups

@@ Abstract.
We find a condition on the acylindrical action of a finitely presented group on a simplicial tree which guarantees that this action will be dominated by an acylindrical action with finitely generated edge stabilisers, and find the first example of an action of a finitely presented group where there is no such dominating action. As a consequence, we show that any finitely presented group that admits a decomposition as an acylindrical graph of (possibly infinitely generated) free groups is virtually compact special, and that every finitely generated subgroup of a one-relator group with an acylindrical Magnus hierarchy is virtually compact special.
@@ 1. Introduction
Let $G$ be a one-relator group. One of the oldest and most effective tools for studying such groups is a Magnus Hierarchy, an iterative HNN-structure on $G$ whose vertex groups are one-relator groups and whose edges are free subgroups generated by subsets of the generators of the vertices (c.f. 3.6). In his paper [undefu], Linton proved that if a Magnus hierarchy for $G$ is acylindrical (c.f. 1.1) then $G$ is hyperbolic and the hierarchy is quasi-convex (c.f. 2.15), a much stronger condition.
Now assume that $G$ has such an acylindrical hierarchy, and let $H$ be a finitely generated subgroup of $G$. By coherence of $G$ [undefq, Theorem 1.1], Gersten’s theorem [undefl, Theorem 5.4], and the fact that one-relator groups have virtual cohomological dimension at most two [undefo, Corollary 9], the group $H$ is hyperbolic, so a natural question to ask is whether $H$ will itself admit a quasi-convex hierarchy. Indeed, $H$ will always admit an acylindrical hierarchy induced by the hierarchy for $G$, but this need not be quasi-convex, the most immediate issue being that we cannot assume that it has finitely generated edge stabilisers — see Section 3.3 for an example of such a $G$ and $H$ where the induced hierarchy on $H$ has infinitely generated edge stabilisers. In this paper we prove that even in such a case, we may construct a quasi-convex hierarchy for $H$.
The definition of an acylindrical action on a tree was first formulated by Sela [undefaf], and later generalised by Weidmann [undefai] to the following.
@@
We say that a group $G$ is acylindrically arboreal if $G$ admits an acylindrical action on a tree that is non-elementary (c.f. Theorem 2.6). Acylindrical arboreality is a special case of acylindrical hyperbolicity, a well studied and powerful generalisation of hyperbolicity that has been of much interest in recent years [undefab, undefad, undefae], and was the first such generalisation to encompass mapping class groups of hyperbolic surfaces ([undeff, undefy], see also [undefad, Section 8]).
A concept strongly related to an acylindrical action on a tree is that of a quasi-convex splitting, an action of a hyperbolic group $G$ on a simplicial tree where the edge stabilisers are all quasi-convex in $G$ (c.f Definition 2.15). A quasi-convex splitting with finitely many orbits of edges will always be acylindrical by a result of Gitik, Mj, Rips and Sageev [undefm, Main Theorem] that quasi-convex subgroups have finite height (c.f. Definition 2.8), meaning that in an action on a tree with quasi-convex edge groups any large enough set of edges will have a finite stabiliser. The converse is more difficult, and is the subject of the following conjecture that can be found in question form in [undefb].
@@
Let $G$ be a hyperbolic group, and $H$ a finitely generated subgroup of $G$ with finite height. Then $H$ is quasi-convex in $G$.
@@
[undefi, Question 1.7] Let $G$ be a hyperbolic group with a non-elementary acylindrical action on a tree. Does $G$ admit a quasi-convex splitting?
One of the main reasons for the significance of quasi-convex splittings is their link to virtual specialness of hyperbolic groups. In particular, in [undefal, Theorem 13.1], Wise proves that if a hyperbolic group admits a quasi-convex splitting with virtually compact special vertex stabilisers then it itself is virtually compact special. Virtual specialness is an important concept in modern geometric group theory, and its definition along with the fact that it is inherited from vertex groups in a quasi-convex splittings was instrumental in proving the Virtual Haken Conjecture, one of the most influential group theoretic results in recent years, see [undef].
@@
    The fundamental groups of closed and orientable hyperbolic $3$-manifolds [undefi, Theorem 1.8]; and
@@
    Free groups, and more generally hyperbolic locally quasi-convex groups such as hyperbolic limit groups [undefaj, Corollary 3.12] and many small cancellation groups [undefz, undefaa].
As stated above, the first obstacle to overcome when attempting to construct quasi-convex splittings from acylindrical ones is that a generic acylindrical splitting need not have finitely generated stabilisers. In [undefi], this problem was solved for hyperbolic $3$-manifold groups using a construction of Stallings, Epstein and Waldhausen which takes a splitting of such a group and outputs a collection of embedded surfaces one may cut along to obtain a new dominating action on a tree. This construction was proved to preserve acylindricity (and indeed promote it to quasi-convexity) in [undefi] using the Subgroup Tameness Theorem for hyperbolic $3$-manifolds (see [undefa, Theorem 5.2], for example).
A similar but weaker method exists in general for finitely presented groups. In particular, there is the construction due to Dunwoody and Sageev known as the Dunwoody–Sageev resolution [undefk, Lemma 2.2] that takes as input an action of a finitely presented group on a tree and outputs a dominating action with finitely generated edge stabilisers that are in some sense geometric, in that they arise as one-dimensional subspaces of a Cayley complex. For a locally quasi-convex group, the actions produced by the Dunwoody–Sageev resolution will always have quasi-convex edge stabilisers, answering Question 1.3 in this case.
The main technical result of this paper exhibits a large class of actions in which acylindricity is preserved by the Dunwoody–Sageev Resolution, and indeed where the resolution produces finite height splittings that dominate the original acylindrical splitting. We say that a group $G$ is locally finite height if every finitely generated subgroup of $G$ has finite height in $G$, and that an action of $G$ on a simplicial tree $T$ is finite height if every edge stabiliser of this action has finite height in $G$. The following is our main theorem.
@@
    If all vertex stabilisers of the action of $G$ on $T$ are locally finite height then $G$ admits a non-elementary finite height acylindrical splitting with finitely generated edge stabilisers that dominates the action of $G$ on $T$. Thus, assuming Conjecture 1.2, if $G$ is hyperbolic then $G$ admits a quasi-convex splitting that dominates the action of $G$ on $T$.
@@
    If there exists an edge stabiliser that is finite height acylindrical in $G$, and this edge stabiliser is locally finite height, then $G$ admits a non-elementary finite height acylindrical splitting with finitely generated edge stabilisers. Furthermore, if every edge stabiliser satisfies this condition then we may assume that the new action dominates the action of $G$ on $T$. Thus, assuming Conjecture 1.2, if $G$ is hyperbolic then $G$ admits a quasi-convex splitting that dominates the action of $G$ on $T$.
We present several applications of this theorem. First, all countable free groups are locally finite height, and thus every finitely presented group that is an acylindrical graph of free groups will admit a finite height splitting with finitely generated edge stabilisers. Indeed, finitely generated free groups are known to be locally quasi-convex by Hall’s Theorem [undefn], and so when dealing with hyperbolic graphs of free groups one may often obtain quasi-convexity without having to appeal to Conjecture 1.2. The following corollary follows immediately from Theorem 1.4(a) and an application of the Quasi-Convex Combination Theorem of Kapovich [undefr, Theorem 1.2], which states that if a group $G$ acts acylindrically on a tree with each vertex stabiliser hyperbolic and each edge stabiliser quasi-convex in its neighbouring vertex stabilisers, then $G$ is hyperbolic each edge stabiliser will be quasi-convex in $G$.
@@
Let $G$ be a finitely presented group, and assume that $G$ acts acylindrically on some tree $T$ such that the vertex stabilisers of this action are free. Then $G$ is hyperbolic and admits a quasi-convex splitting with free vertex groups, and so $G$ is virtually compact special by [undefal, Theorem 13.1].
Our methods also apply to subgroups of hyperbolic groups with quasi-convex splittings over locally finite height groups. The following corollary allows us to obtain virtual compact specialness for many groups without appealing to Conjecture 1.2.
@@
Let $G$ be a hyperbolic group, and let $H$ be a hyperbolic subgroup of $G$.
  1. (a)
    If $G$ admits an action on a tree $T$ that is a quasi-convex splitting over a locally quasi-convex group, and $H$ the induced action of $H$ on $T$ is non-trivial, then $H$ admits a non-trivial quasi-convex splitting that dominates the minimal subtree of the induced action of $H$ on $T$.
  2. (b)
    If $G$ admits a quasi-convex hierarchy with locally quasi-convex edge stabilisers then $H$ admits a quasi-convex hierarchy, and therefore $H$ is virtually compact special by [undefal, Theorem 13.1].
In the case where $G$ has virtual cohomological dimension $2$, and so we can apply [undefl, Theorem 5.4], we need only ask for $H$ to be finitely presented. For example, the following settles our opening question.
Corollary 1.7.
Let $H$ be a finitely generated subgroup of a one-relator group $G$ such that $G$ admits an acylindrical Magnus hierarchy $G_{0}\hookrightarrow G_{1}\hookrightarrow\cdots\hookrightarrow G_{n}\cong G$. Then $H$ admits a a quasi-convex hierarchy, and so is virtually compact special by [undefal, Theorem 13.1]. In particular the group $H$ is acylindrically arboreal if and only if it admits a non-trivial a quasi-convex splitting.
Furthermore, $H$ admits a quasi-convex hierarchy and in particular is virtually compact special by [undefal, Theorem 13.1].
This corollary will apply to subgroups of many two-generator one-relator groups, such as the one-relator group given in Example 3.9, and subgroups of all one-relator groups with negative immersions [undefu, Corollary 6.9].
@@
A positive answer to Question 1.8 would have strong consequences. In particular, again using [undefal, Theorem 13.1], a positive answer would imply that every hyperbolic group that admits an acylindrical hierarchy in fact admits a quasi-convex hierarchy, and therefore is virtually compact special with a proof analogous to that of Corollary 1.7. This would provide a significantly weaker condition for virtual specialness of a hyperbolic group than we currently have at our disposal.
We finish by illustrating the issues that arise when attempting to relax the hypotheses of Theorem 1.4. In Section 3.4 we will show that the Dunwoody–Sageev resolution cannot be assumed to preserve the acylindricity of a general action with the following proposition.
@@
In particular, the Dunwoody–Sageev resolution will not preserve acylindricity in general.
Somewhat unfortunately, the group constructed does admit an acylindrical action with finitely generated edge stabilisers. To counter this, we ask the following.
Question 1.10.
@@ 1.1. Acknowledgements
This work was completed while the author was a PhD student at the University of Cambridge, supervised by Jack Button, and the author is very grateful to his supervisor for all of his help. The author would also like to thank Henry Wilton for several very helpful discussions, and Daniel Groves for his helpful suggestions. Finally, the author is very grateful to an anonymous reviewer, whose careful reading and helpful suggestions allowed for a much shortened proof of our main theorem, and for Corollaries 1.6 and 1.7 to be proved independently from Conjecture 1.2.
Financially, support from the Cambridge Trust Basil Howard Research Graduate Studentship is gratefully acknowledged.
@@ 2. Preliminaries - Acylindrical Hyperbolicity and Acylindrical Arboreality
We will assume the reader has some familiarity with Bass–Serre theory, and for a more detailed discussion we refer to [undefag, undefj]. We define a graph of groups $(\Gamma,\mathfrak{G})$ to be a connected directed graph $\Gamma$ that may not be finite or simple along with $\mathfrak{G}$, which is the following data:
@@
We define the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ of $(\Gamma,\mathfrak{G})$ as in [undefag, Section I.5.1].
We will use a slight abuse of notation to consider each vertex group $G_{v}$ as a subgroup of the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ along the natural inclusion. Similarly, we will consider each edge group $G_{e}$ to be the subgroup of the fundamental group given by the image of $d_{0}(G_{e})$ in the vertex group $G_{i(e)}$. We call a graph of groups trivial if there exists some $v\in V(\Gamma)$ such that $G_{v}=\pi_{1}(\Gamma,\mathfrak{G})$, or non-trivial otherwise. We say that a graph of groups $(\Gamma,\mathfrak{G})$ is a graph of groups decomposition or splitting of a group $G$ if the fundamental group $\pi_{1}(\Gamma,\mathfrak{G})$ is isomorphic to $G$, and identify $G$ with $\pi_{1}(\Gamma,\mathfrak{G})$. We denote by $T(\Gamma,\mathfrak{G})$ the Bass–Serre tree associated to the splitting, on which $G$ acts naturally by isometry with respect to the edge metric and without inversion as in [undefag, Section I.5.3].
From now on we will assume that all trees are simplicial, and all actions on trees are by simplicial isometry and without inversion, so that every action on a tree is associated to a quotient graph of groups as in [undefag, Section I.5], whose edge and vertex groups correspond to conjugacy classes of stabilisers of the original group action. We say that an action of a group $G$ on a tree $T$ is minimal if there exists no proper subtree of $T$ that is fixed setwise by the action of $G$. We have the following lemma, which will allow us to assume that the actions we consider are minimal in most cases.
@@
[undefj, Lemma I.4.13] Let $G$ be a finitely generated group acting on a tree $T$. Then there exists a unique minimal subtree $T_{m}\subseteq T$, and the induced action of $G$ on $T_{m}$ has finitely many orbits of edges.
The following lemma of Bieri will be essential in dealing with finite generation of stabilisers of actions on trees.
@@
[undefc, Proposition 2.13(a) and (b)] Let $G$ be a finitely generated group acting on a tree $T$ with finitely many orbits of edges. If every edge stabiliser of this action is finitely generated, then so is every vertex stabiliser.
@@
    For a more complicated example, a right angled Artin group is acylindrically arboreal if and only if the underlying graph has graph theoretical diameter at least three [undefi, Corollary 3.12].
@@
    If $G$ is the fundamental group of a closed non-geometric $3$-manifold, then Wilton and Zalesskii prove that $G$ acts acylindrically on the tree corresponding to its JSJ-decomposition [undefak, Lemma 2.4]. In the hyperbolic case, the fundamental group $G$ of a finite volume orientable hyperbolic $3$-manifold will be acylindrically arboreal if and only if the manifold contains an embedded $2$-sided closed incompressible surface whose fundamental group includes into $G$ as a geometrically finite subgroup [undefi, Theorem 1.6].
By a result first explicitly proved in [undefi, Theorem 2.17], but essentially due to Minasyan and Osin [undefab, Lemma 4.2], an action on a simplicial tree is non-elementary acylindrical as in Definition 2.4 if and only if it is non-elementary acylindrical in the sense of Bowditch [undeff]. Thus, if a group is acylindrically arboreal then it is acylindrically hyperbolic, and we have the following powerful classification theorem.
@@
[undefad, Theorem 1.1] Let $G$ be a group acting acylindrically on a tree $T$. Then $G$ satisfies exactly one of the following conditions.
@@
[undefad, Corollary 1.5] Let $G$ be an acylindrically arboreal group. Then every infinite normal subgroup of $G$ is acylindrically arboreal. In particular, such a subgroup cannot be abelian.
2.1. finite height and Quasi-Convex Splittings
As in Question 1.3, we are concerned with whether we can use acylindrical splittings to build splittings with stronger properties. In particular, we are concerned with finite height and quasi-convex splittings of hyperbolic groups.
Intuitively, a finite height splitting is an action on a tree where a set of edges containing enough elements has a finite stabiliser. It is important to note that, unlike acylindricity, this definition does not require that these stabilisers are uniformly bounded, only that they are finite. More formally, we define finite height as follows.
@@
Let $G$ be a group, and $H\leq G$ a subgroup. For $g_{1},g_{2}\in G$, we say that the conjugates $g_{1}Hg_{1}^{-1}$ and $g_{2}Hg_{2}^{-1}$ are essentially distinct if $g_{1}H\neq g_{2}H$.
For some natural number $n$, we say that $H$ has height at most n in $G$ if the intersection of any collection of at least $n$ essentially distinct conjugates of $H$ in $G$ is finite. We define the height of $H$, $h_{G}(H)$ to be the smallest $m$ such that $H$ has height at most $m$ in $G$, or to be infinite if no such $m$ exists.
@@
Some of the simplest examples of finite height subgroups are finite index subgroups. If $G$ is a group and $H\leq G$ has finite index then $H$ will have height at most $[G:H]+1$, owing to the fact that no collection of $[G:H]+1$ left cosets of $H$ in $G$ can be disjoint.
We say that the action of a group $G$ on a tree $T$ is a finite height splitting if for all edges $e\in E(T)$, the $G$-stabiliser $G_{e}$ of $e$ is finite height in $G$, i.e $h_{G}(G_{e})$ is finite. As above, some care should be taken when comparing acylindricity to finite height splittings owing to the different restrictions on the finite stabilisers. In particular, while in an acylindrical action stabilisers of long enough paths must be uniformly bounded, in a finite height splitting we only require that stabilisers of long enough sets are finite. The following example demonstrates this difference.
@@
The group $G$ acts on $T$ via the left coset action on the labels of its vertices, and the edge stabilisers of this action are all isomorphic to finite groups, so naturally have height $1$. This action is therefore finite height, but will not be acylindrical, as for all $k$ we can find a path of length $k$ whose stabiliser is contains $2^{i}$ elements for any $i$ by taking the path that starts at the vertex labelled $(\mathbb{Z}/2^{i}\mathbb{Z})$ and finishing at the vertex labelled $(\mathbb{Z}/2^{i+k}\mathbb{Z})$.